OpenStudy (anonymous):

LOOKS EASY! What is the next term in the following constant-velocity sequence? 10 cm, 16 cm, 22 cm,

OpenStudy (anonymous):

so constant velocty is 6 cm per step?

OpenStudy (anonymous):

10/2=5 16/2=8 22/2=11 Answer=28

OpenStudy (anonymous):

i didn't do that, i added 6 each time and got 28 but ok